TWO-COLOR RESONANT FOUR-WAVE MIXING SPECTROSCOPY OF HIGHLY PREDISSOCIATED LEVELS IN THE A^{2}A_{1} STATE OF CH_{3}S

C.-P. LIU, Department of Chemistry, National Tsing Hua University, Hsinchu 30013, Taiwan; \underline{S. A. REID}, Department of Chemistry, Marquette University, Milwaukee, WI 53233; Y.-P. LEE, Department of Applied Chemistry, National Chiao Tung University, Hsinchu 30010 and Institute for Atomic and Molecular Sciences, Academia Sinica, Taipei 106, Taiwan.

We report results of two-color resonant four-wave mixing experiments on highly predissociated levels of the methylthio radical, CH_{3}S in its first excited electronic state. The radical was generated by 248 nm photolysis of dimethyl disulfide, and the spectra measured in a hole-burning scheme in which the probe laser excited specific rotational transitions in band 3^{3}. The spectral simplification afforded by the two-color method allowed accurate determination of line positions and homogeneous linewidths, which are reported for the C-S stretching states 3^{n} with n=3-7 and combination states 1^{1}3^{n} (n=0-2), 2^{1}3^{n} (n=3-6), and 1^{1}2^{1}3^{n} (n=0,1). The spectra show pronounced mode specificity, as the homogeneous linewidth of levels with similar energies varies by up to two orders of magnitude. The derived vibrational parameters of the A^{2}A_{1} state are in satisfactory agreement with ab initio predictions.